OpenStudy (anonymous):

A box has sides with these measures: length 10 cm, width 8 cm, height h cm. If the volume is 400 cm3 , find the measure of the height. You may use the equation V = lwh as a beginning. A. 0.2 cm B. 4 cm C. 0.5 cm D. 5 cm

sam (.sam.):

V=L*W*H 400 = 10*8*H 400 = 80H (400/80)=(80H)/80 5=H D